OpenStudy (anonymous):

use the subsitution to solve the system of linear equations 3x-y=8 7x-2y=21

OpenStudy (anonymous):

ok do you understand the term "Substitution?"

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok first thing rearrange the first equation to the general linear form y=mx + c 1) 3x - y = 8 3x -8 = y y=3x-8 Now we want to substitute y=3x-8 into the second equation given which is 7x-2y=21 7x - 2y = 21 7x - 2(3x-8) = 21 7x -6x + 16 = 21 x + 16 = 21 x = 21-16 x = .......... (I'm sure u can do this)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

yes that is correct.

OpenStudy (anonymous):

but do u understand the concept?

OpenStudy (anonymous):

Thankk Youu !

OpenStudy (anonymous):

Yess Andd Will 5 Be Myy Ordered Pair ?

OpenStudy (anonymous):

nope, that is just your x-value for your ordered pair.

OpenStudy (anonymous):

there is one more step, just substitute x=5 into either of the original equation to find your y-value. currently you have (5,?) we need to find the ? which is your y-value of the ordered pair.

OpenStudy (anonymous):

Okai And How Do I DO That

OpenStudy (anonymous):

simply substitute x=5 into either of the original equations which are 3x-y=8 or 7x-2y=21

OpenStudy (anonymous):

Okaii Im Stuckk Can You Please Help Me Out

OpenStudy (anonymous):

ok, we figured x=5 already and we choose one of the original equations, lets choose 3x-y=8 remember substitute means "replace" so we replace x=5 into that equation we choose. 3x-y=8 3(5) - y = 8 15 - y = 8 y = ....... (you can do this)

OpenStudy (anonymous):

7 ?

OpenStudy (anonymous):

yes that is correct, now tell me the final ordered pair which is your solution/answer.

OpenStudy (anonymous):

(4,7)

OpenStudy (anonymous):

no.... (5,7)

OpenStudy (anonymous):

remember we found x=5 at first, how can it become a 4?

OpenStudy (anonymous):

Oh Yes Im So Sorry i Had Forgot AbOut This ! Its Confusing !

OpenStudy (anonymous):

No Can You Please Help Me With This One To use the method of substitution to solve linear equations 7a=3-b and 7a=7-b

OpenStudy (anonymous):

use the eliminatiuon method to solve the system of equations a+8b=1 a+4b=-3

OpenStudy (anonymous):

And What Will The A Be?

OpenStudy (anonymous):

u have to solve it and tell me.

OpenStudy (anonymous):

Will I Use The A As A 1 ?